OpenStudy (dls):

If y(x) satisfies the differential equation y'-ytanx=2xsecx and y(0)=0 then?

OpenStudy (dls):

A) y(pi/4)=pi^2/8 sqrt(2) B)y'(pi/4)=pi^2/18 C)y(pi/3)=pi^2/9 D)y'(pi/3)=4pi/3+2pi^2/3sqrt(3)

OpenStudy (dls):

Attempt: I guess its a linear differential equation of type \[\LARGE \frac{dy}{dx}+PY=Q\] Where, P= -tanx And integrating factor, \[\LARGE I.F=e^{\int\limits pdx}=cosx\] After solving a bit,: \[\LARGE Y \cos x=x^2+c\] According to given conditions,\(\large c=y\) \[\LARGE Y(cosx-1)=x^2\] \[\LARGE y=\frac{x^2}{cosx-1}\] This is my equation,which doesn't prove any option right..

OpenStudy (primeralph):

The "after solving a bit" part might be where the problem is.

OpenStudy (primeralph):

c=0 not y.

OpenStudy (primeralph):